OpenStudy (dancing_stars23):

How many different 12-member juries can be made from 150 people? okay so I know the formula is 150C12 = 150!/12!138! but when I put this into a calculator, it says error because it is too high for a number to compute.... so what do I do?

Nnesha (nnesha):

\[\frac{ 150! }{ 138! 12!} = \frac{ 150 \times 149 \times 148..... }{ 12! 138!}\] 150! subtract one until you get 138 at the denominator the you can cancel out top 138! by 138! at the denominator

OpenStudy (dancing_stars23):

1.72420656⋅10^17 ?

OpenStudy (dancing_stars23):

@Nnesha

OpenStudy (dancing_stars23):

well okay so I was able to ask my brother and he said I can go like this since they are really big numbers....

OpenStudy (dancing_stars23):

\[(150*149*148*147*146*145*144*143*142*141*140*139)/12!\] and you will cancel out the 138.

OpenStudy (dancing_stars23):

he's in college and I wasn't sure if I was going to get any help on here, and so I asked. and then I wanted to make sure my answer was correct

Nnesha (nnesha):

yep like that \[(150*149*148*147*146*145*144*143*142*141*140*139 \times \cancel{138!})/12! \cancel{138!}\] yep looks right

OpenStudy (dancing_stars23):

okay thanks!

Nnesha (nnesha):

np :-)